OpenStudy (anonymous):

log {2^(x+2) - 4^(x) } of base 1/3 >= (-2) can any one tell the method to solve

OpenStudy (anonymous):

heyy its base is 10 na????

OpenStudy (lgbasallote):

\(\Huge \log_{\frac 13} 2^{(x+2)} - 4^x \ge -2\) ^right?

OpenStudy (anonymous):

yup the format is the one lgbasallote typed

OpenStudy (dumbcow):

well you could get rid of log by changing to exponential form \[2^{x+2} -4^{x} \ge (1/3)^{-2}\] simplify so you have same base \[4*2^{x} -2^{2x} \ge 9\] let u = 2^x it becomes a quadratic \[u^{2}-4u+9 \le 0\] use quadratic formula ...

OpenStudy (dumbcow):

finally once you have the 2 roots , u1 and u2 \[u_{1} \le 2^{x} \le u_{2}\]